quickstudent:

A Ferris Wheel at the amusement park is in the shape of a regular 18-gon. What is the measure of each interior angle of the polygon? 2880° 20° 22.5° 160°

EvL:

measure of each interior-angle in an n-polygon: \(\Large \frac{(n-2)\times 180}{n}\), where n=18

quickstudent:

160